OpenStudy (anonymous):

Which operation would you use on the exponents to simplify the expression the quantity nine to the fourth raised to the negative second power? Multiplication Division Addition Subtraction

OpenStudy (anonymous):

flutter

OpenStudy (anonymous):

feather

OpenStudy (anonymous):

pellet

zepdrix (zepdrix):

whut? lol

OpenStudy (anonymous):

they are changing my curse words to weird words

zepdrix (zepdrix):

Oh there's a filter? LOL I didn't even know that! :D

zepdrix (zepdrix):

Here is how we would apply our rule for exponents like this: \[\Large \left(9^4\right)^{-2}\quad=\quad 9^{4\cdot-2}\] So what operation is being performed?